Please help me solve this equation: (3n+6)(6n-3)=0
Thanks for your help.

Found 2 solutions by algebrahouse.com, robertb:
Answer by algebrahouse.com(1659)

"Please help me solve this equation: (3n+6)(6n-3)=0
Thanks for your help."

If you have two terms multiplied together and equal to 0,
then either one of them could be 0.

To find all the possibilities for n, set each term (factor) equal to 0.

(3n + 6)(6n - 3) = 0
3n + 6 = 0 or 6n - 3 = 0 {set each factor equal to 0}
3n = -6 or 6n = 3 {solving both equations for n}
n = -2 or n = 1/2
www.algebrahouse.com

Answer by robertb(5830)

3n + 6 = 0 ==> n = -2
6n - 3 = 0 ==> n = 1/2